OpenStudy (anonymous):

Evaluate b - 2a - c for a = -7, b =3, and c = 7 A. 24 B. -18 C. 3 D. 10

OpenStudy (nurali):

b - 2a - c a = -7, b =3, and c = 7 3-2*-7-7 3+14-7 3+7 10

OpenStudy (anonymous):

Thanks, was wrong but gave you a medal anyways!

OpenStudy (anonymous):

3 - 2(-7)- 7 3 -(-14) -7 =3 + 14 -7 =10